# SelfCite: Self-Supervised Alignment for Context Attribution in Large Language Models

### v1

@@ Abstract
We introduce SelfCite, a novel self-supervised approach that aligns LLMs to generate high-quality, fine-grained, sentence-level citations for the statements in their generated responses. Instead of only relying on costly and labor-intensive annotations, SelfCite leverages a reward signal provided by the LLM itself through context ablation: If a citation is necessary, removing the cited text from the context should prevent the same response; if sufficient, retaining the cited text alone should preserve the same response. This reward can guide the inference-time best-of-N sampling strategy to improve citation quality significantly, as well as be used in preference optimization to directly fine-tune the models for generating better citations. The effectiveness of SelfCite is demonstrated by increasing citation F1 up to 5.3 points on the LongBench-Cite benchmark across five long-form question answering tasks.
[…]
Yung-Sung Chuang , Shang-Wen Li , Scott Yih
@@ 1 Introduction
Although completely eliminating hallucinations remains difficult, existing approaches have sought to enhance the reliability of LLMs by providing context attributions–commonly referred to as citations–which are fine-grained references to relevant evidences from the context, alongside generated responses for user verification (Menick et al., 2022; Slobodkin et al., 2024; Zhang et al., 2024). While they have shown promise in generating citations, an outstanding challenge is their reliance on annotated data either from human (Menick et al., 2022; Slobodkin et al., 2024) or costly proprietary APIs (Zhang et al., 2024; Huang et al., 2024) to train models to generate citations. Collecting annotations can be time-consuming and costly, especially with long-context documents.
@@
Building on this intuition, we design a reward that can be cheaply computed by the LLM itself, composed by probability drop and probability hold in context ablation. By integrating this reward function into a best-of-N sampling strategy, SelfCite achieves substantial improvements in citation quality. Furthermore, we employ this reward for preference optimization using SimPO (Meng et al., 2024), which not only maintains these improvements but also eliminates the need for best-of-N sampling. We outperform the previous state of the art on the LongBench-Cite benchmark (Zhang et al., 2024) by up to 5.3 points in F1 scores, and showing a promising direction to bootstrap the citation quality from LLMs via self-rewarding.
@@ 2.1 Problem Formulation
Consider employing an autoregressive language model (LM) to generate a response to a specific query given a context of relevant information. Specifically, given an LM $p_{\text{LM}}$, let $p_{\text{LM}}(t_{i}\mid t_{1},\ldots,t_{i-1})$ denote its output distribution over the next token $t_{i}$ based on a sequence of preceding tokens $t_{1},\ldots,t_{i-1}$. Next, let $C$ represent the context of relevant information. This context is partitioned into $|C|$ sentences: $c_{1},c_{2},\dots,c_{|C|}$. Each sentence $c_{j}$ is prepended with a unique identifier (e.g. sentence index $j$) as a way for the model to reference the sentence when generating citations. The context $C$ is followed by a query $Q$, a question or instruction for the model. A response $R$ is then sampled from the LM $p_{\text{LM}}$.
@@ 2.2 Self-Supervised Reward via Context Ablation
$\text{Reward}(e_{i})=\log p_{\text{LM}}\left(r_{i}|E_{i}\right)-\log p_{\text{LM}}\left(r_{i}|C\setminus E_{i}\right).$ (1)
@@ 2.3 Best-of-N Sampling
To leverage the self-supervised reward computed via context ablation, we employ a best-of-N sampling strategy, which is a common way to test the effectiveness of a reward design (Gao et al., 2023a; Lightman et al., 2024) as a performance oracle without any confounders from training. After generating the full response, we locate the position where the citation tags <cite>...</cite> are generated. Within the citation tags, we sample $N$ candidate citation sequences and select the citation set that maximizes the combined reward metric, Eq. (1). The corresponding procedure is shown in Algorithm 1. After obtaining the selected citations $\{e_{1}^{*},\dots,e_{S}^{*}\}$, we replace the original citation sequence $e_{i}$ with the optimal citation set $e_{i}^{*}$ for the response statement $r_{i}$, while keeping the response statements $\{r_{1},\dots,r_{S}\}$ unchanged. This process is repeated for each statement in the response $R$ to obtain the final, citation-improved output $R^{*}=\{r_{1},e_{1}^{*},\dots,r_{S},e_{S}^{*}\}$. To prevent the model from citating too many sentences, we exclude the BoN candidates that cites more than 384 tokens in total, unless they are all from a single sentence.
@@
0:  LM $p_{\text{LM}}$, context $C$, query $Q$, response $R$, number of candidates $N$
@@
        $\mathrm{Reward}(e_{i}^{(k)})$ $=\log p_{\text{LM}}\bigl{(}r_{i}\mid E_{i}^{(n)}\bigr{)}-\log p_{\text{LM}}\bigl{(}r_{i}\mid C\setminus E_{i}^{(n)}\bigr{)}$
@@
     $e_{i}^{*}=\arg\max_{k}\;\mathrm{reward}\bigl{(}e_{i}^{(k)}\bigr{)}$
@@ 2.4 Preference Optimization
Best-of-N sampling is a straightforward way to obtain better citations, but at the additional inference time of generating candidates and reranking. Thus, we try to internalize the ability of generating better citations back to the LM itself.
Given documents and queries, we can prompt the LM to generate the responses along with the citations $R=\{r_{1},e_{1},...,r_{S},e_{S}\}$. By further applying best-of-N sampling, we can obtain new responses of the same statements but with better citations $R^{*}=\{r_{1},e^{*}_{1},...,r_{S},e^{*}_{S}\}$. Such preference data can be used in direct preference optimization (DPO) (Rafailov et al., 2024) to align the model based on the preference between the original outputs and improved outputs. DPO typically requires more memory usage than SFT due to the need of a reference model. Also, optimizing with preference data pairs inherently makes the per-GPU batch size to be at least 2, limiting the maximum context length that can be used. To address these challenges, we use SimPO (Meng et al., 2024), a variant of DPO that does not require a reference model, freeing up memory for long-context fine-tuning. We further apply Liger-Kernel (Hsu et al., 2024), a collection of efficient Triton kernels, to optimize memory usage and scale the context length to 25.6K, as detailed in Appendix 7. Through this self-supervised alignment process, which does not require ground-truth answers or human annotations, the model learns to generate more accurate and contextually grounded citations on its own.
@@ 3.1 Model Details
We use the Llama-3.1-8B model (Dubey et al., 2024) fine-tuned on LongCite-45K SFT data, namely the LongCite-8B model (Zhang et al., 2024) as the start point for both best-of-N sampling and preference optimization. We adopt the same text segmentation strategy from Zhang et al. (2024): each document is split into individual sentences using NLTK (Bird, 2006) and Chinese punctuations, and each sentence is prepended with a unique identifier in <C{$i$}> format. These identifiers serve as the citation indices, enabling the model to cite relevant context right after the statements with the format of <statement> {content ...} <cite>[$i_{1}-i_{2}$][$i_{3}-i_{4}$]...</cite></statement>. This format allows the model to cite a single sentence (e.g. $i_{1}=i_{2}$) or a span (e.g. $i_{1}<i_{2}$) efficiently within several tokens. The responses are generated via top-p sampling (Holtzman et al., 2020) with p=0.7 and temperature=0.95. We set p=0.9 and temperature=1.2 when doing best-of-N sampling for the citation strings to increase the diversity. We set N=10 in all the experiments considering the limited diversity in citations. [^1] After deduplicating repeated citation candidates, on average there are only 4.8 candidates left per statement in the BoN experiment on LongBench-Cite, with a standard deviation of 3.2.
@@ 3.2 Preference Optimization
Since best-of-N responses tend to have slightly longer citations, directly fine-tuning on them can lead the model to adopt a shortcut—generating longer citations instead of improving citation quality. To prevent this, we introduce length balancing: if an original response has a shorter citation length than the best-of-N response, we insert random citations from nearby sentences. This encourages the model to focus on where to cite rather than simply citing more. Details are provided in Appendix 9, with an ablation study in Section 4.2.
3.3 Evaluation
Benchmark.
Baselines.
SelfCite is compared with these baselines.
  • Contributive context attribution: Contributive context attribution seeks to directly identify the parts of the context that cause the model to generate a particular statement. We consider ContextCite (Cohen-Wang et al., 2024), a contributive context attribution method that performs several random context ablations to model the effect of ablating different parts of the context on a generated statement. We apply ContextCite with 32 and 256 times of random context ablations, with the details described in Appendix 8.
  • Fine-tuned models: LongCite-8B and 9B released by Zhang et al. (2024), trained on LongCite-45K, fine-tuned from Llama-3.1-8B (Dubey et al., 2024) and GLM-4-9B (GLM et al., 2024), respectively.
Table 1: Citation recall (R), citation precision (P), citation F1 (F1), and citation length evaluated on LongBench-Cite benchmark. The best of our results are bolded. The best of previous state of the art are underlined. indicates the results taken from Zhang et al. (2024). Our repro. means our reproduced results.
@@ 3.4 Main Results
Table 1 presents our main results. Our best-of-N sampling (BoN) consistently improves both citation recall and citation precision across tasks, increasing the overall F1 score from 73.8 to 77.5. Using SimPO to internalize BoN’s gains—eliminating the need for BoN sampling—achieves a similar improvement, with an F1 of 77.9. Applying BoN again to the SimPO fine-tuned model further boosts F1 by 5.3 points to 79.1, the highest across the datasets, suggesting room for further gains. Our results surpass LongCite-8B/9B at similar citation lengths and outperform proprietary model prompting while producing shorter citations. Compared to ContextCite, which also relies on context ablation, our approach is significantly better. A key reason is that ContextCite estimates sentence importance from scratch using linear regression, while we rerank existing LLM-generated citation candidates, leading to more efficient and accurate citation quality estimation. Additionally, we evaluate the latest released Claude Citations API, as shown in Appendix 10 that SelfCite achieves strong results very close to this commercial-level API that specialized for providing citations.
@@
In our main experiment, we start from the Llama-3.1-8B model fine-tuned on the LongCite-45K SFT data, which effectively kick-starts its ability to generate structured citations for best-of-N sampling. The subsequent SimPO alignment stage is entirely self-supervised. We are also curious if it is possible to start from a fully self-supervised SFT model and then apply our self-supervised alignment after that. To begin with, we automatically generate 11K citation SFT data using ContextCite (see Appendix 8 for details) to replace the LongCite-45K annotations in the training data, as shown in the results at the bottom of Table 1. We can see that SFT on ContextCite can achieve decent initial results (65.7 F1) but still far from LongCite-8B (73.8 F1). BoN helps improving F1 to 67.3. After SimPO training, it achieves 69.9 F1, and additionally applying BoN can boost its F1 by 5.8 to 71.5, significantly closing the gap to LongCite-8B, showing our alignment method not only improve the supervised models, but also enhance the models purely trained from self-supervision.
@@
For best-of-N sampling, only the citation parts are modified, so the responses it generates to answer the questions are the same as those of the original LongCite-8B model, maintaining the same correctness. For the SimPO finetuned models, we test their answer correctness by the evaluation in Zhang et al. (2024), which contains two settings: answering with/without citations. If answering with citations, the citation parts will be removed when evaluating the answer correctness. The results in Table 2 show that the SimPO finetuning does not change the correctness of the LongCite-8B model much. The correctness is similar to LongSFT-8B/9B (Zhang et al., 2024), which are ablation baselines finetuned on LongCite-45k QA pairs but without the citation parts. When finetuning Llama-3.1-8B-Instruct with ContextCite SFT data, the answer correctness slightly degrades. It is probably caused by the fact that we do not mix any instruction following dataset in the SFT stage as Zhang et al. (2024) does. However, the further SimPO step does not change the answer correctness significantly, while lead to better citations for user verification. Since the SFT stage is not our main method here, we leave the better data mixture strategy in SFT stage in the future work.
@@ 4.1 Ablation Study on Rewards
To better understand our final reward design, we explore various reward strategies in the BoN sampling process. Here, all BoN candidates are pre-generated and fixed, the reward is the only factor affecting results. Table 3 presents our ablation results on HotpotQA, while citation lengths are computed across all LongBench-Cite datasets for direct comparison with Table 1. We evaluate four alternative reward designs. BoN by LM log prob re-ranks candidates simply by the probability of the citation string, <cite>[$i_{1}-i_{2}$][$i_{3}-i_{4}$]...</cite>, which is similar to the effects of beam search. We observe that this strategy slightly boosts recall while reducing precision, resulting in a minor reduction in F1. BoN by max citation length always selects the candidates with the longest citations, i.e. citing the greatest number of sentences. Although it improves recall, it significantly reduces precision from 77.9 to 73.6 and inflates the citation length from 83.5 to 139.8. By contrast, both BoN by Prob-Drop and BoN by Prob-Hold improve recall without sacrificing precision. Finally, by combining both Prob-Drop and Prob-Hold into our final SelfCite reward, we achieve the best outcome, increasing both recall and precision and a 4-point improvement in F1.

### v2

@@ Abstract
We introduce SelfCite, a novel self-supervised approach that aligns LLMs to generate high-quality, fine-grained, sentence-level citations for the statements in their generated responses. Instead of only relying on costly and labor-intensive annotations, SelfCite leverages a reward signal provided by the LLM itself through context ablation: If a citation is necessary, removing the cited text from the context should prevent the same response; if sufficient, retaining the cited text alone should preserve the same response. This reward can guide the inference-time best-of-N sampling strategy to improve citation quality significantly, as well as be used in preference optimization to directly fine-tune the models for generating better citations. The effectiveness of SelfCite is demonstrated by increasing citation F1 up to 5.3 points on the LongBench-Cite benchmark across five long-form question answering tasks. The source code is available at https://github.com/facebookresearch/SelfCite.
Machine Learning, ICML
@@ 1 Introduction
Although completely eliminating hallucinations remains difficult, existing approaches have sought to enhance the reliability of LLMs by providing context attributions–commonly referred to as citations–which are fine-grained references to relevant evidences from the context, alongside generated responses for user verification (Menick et al., 2022; Slobodkin et al., 2024; Zhang et al., 2024). While they have shown promise in generating citations, an outstanding challenge is their reliance on annotated data either from human (Menick et al., 2022; Slobodkin et al., 2024) or costly proprietary APIs (Zhang et al., 2024) to train models to generate citations. Collecting annotations can be time-consuming or costly, especially with long-context documents.
@@
Building on this intuition, we design a reward that can be cheaply computed by the LLM itself, composed by probability drop and probability hold in context ablation. By integrating this reward function into a best-of-N sampling strategy, SelfCite achieves substantial improvements in citation quality. Furthermore, we employ this reward for preference optimization using SimPO (Meng et al., 2024), which not only maintains these improvements but also eliminates the need for the computationally expensive best-of-N sampling. We outperform the previous state of the art on the LongBench-Cite benchmark (Zhang et al., 2024) by up to 5.3 points in F1 scores, and showing a promising direction to bootstrap the citation quality from LLMs via self-rewarding.
@@ 2.1 Problem Formulation
Consider employing an autoregressive language model (LM) to generate a response to a specific query given a context of relevant information. Specifically, given an LM $p_{\text{LM}}$, let $p_{\text{LM}}(t_{i}\mid t_{1},\ldots,t_{i-1})$ denote its output distribution over the next token $t_{i}$ based on a sequence of preceding tokens $t_{1},\ldots,t_{i-1}$. Next, let $C$ represent the context of relevant information. This context is partitioned into $|C|$ sentences: $c_{1},c_{2},\dots,c_{|C|}$. Each sentence $c_{j}$ is prepended with a unique identifier (e.g., sentence index $j$) as a way for the model to reference the sentence when generating citations. The context $C$ is followed by a query $Q$, a question or instruction for the model. A response $R$ is then sampled from the model $p_{\text{LM}}$.
@@ 2.2 Self-Supervised Reward via Context Ablation
$\text{Reward}(e_{i})=\text{Prob-Drop}(e_{i})+\text{Prob-Hold}(e_{i})$
$=\log p_{\text{LM}}\left(r_{i}|E_{i}\right)-\log p_{\text{LM}}\left(r_{i}|C\setminus E_{i}\right).$ (1)
@@ 2.3 Best-of-N Sampling
To leverage the self-supervised reward computed via context ablation, we employ a best-of-N sampling strategy, which is a common way to test the effectiveness of a reward design (Gao et al., 2023a; Lightman et al., 2024) as a performance oracle without any confounders from training. For convenience, we first generate the full response $R=\{r_{1},e_{1},\dots,r_{S},e_{S}\}$ which includes a set of statements ($r_{i}$) paired with citations ($e_{i}$), and then locate the position of $e_{i}$, i.e., where the citation tags <cite>...</cite> are generated. Within the citation tags of $e_{i}$, we re-sample $N$ candidate citation sequences ($e_{i}^{(1)},\dots,e_{i}^{(N)}$), by making the model to continue the generation from $\{C,Q,r_{1},e_{1},\dots,r_{i}\}$, and then select the best citation ($e_{i}^{*}$) that maximizes the combined reward metric, Eq. (2.2). The corresponding procedure is shown in Algorithm 1. After obtaining all the selected citations $\{e_{1}^{*},\dots,e_{S}^{*}\}$, we replace the original citation sequence $e_{i}$ with the optimal citation $e_{i}^{*}$ for each response statement $r_{i}$, while keeping the response statements $\{r_{1},\dots,r_{S}\}$ unchanged. This process is repeated for each statement in the response $R$ to obtain the final, citation-improved output $R^{*}=\{r_{1},e_{1}^{*},\dots,r_{S},e_{S}^{*}\}$. To prevent the model from citing too many sentences, we exclude the candidate $e_{i}$ if the cited text ($E_{i}$) is longer than $L_{\text{max}}=384$ tokens in total, unless $E_{i}$ are all from a single long sentence.
@@
[…]
@@
        if $T(E_{i}^{(k)})<=L_{\text{max}}\text{ or }\text{\#}(e_{i}^{(k)})=1$ then
           $\mathrm{Reward}(k)$ $=\log p_{\text{LM}}\bigl{(}r_{i}|E_{i}^{(k)}\bigr{)}-\log p_{\text{LM}}\bigl{(}r_{i}|C\setminus E_{i}^{(k)}\bigr{)}$
        end if
@@
     $k^{*}=\arg\max_{k}\;\mathrm{Reward}(k)$
     $e_{i}^{*}=e_{i}^{(k^{*})}$
@@ 2.4 Preference Optimization
Best-of-N sampling is a straightforward way to obtain better citations, but at the additional inference cost of generating candidates and reranking. Thus, we try to internalize the ability of generating better citations back to the LM itself.
Given documents and queries, we can prompt the LM to generate the responses along with the citations $R=\{r_{1},e_{1},...,r_{S},e_{S}\}$. By further applying best-of-N sampling, we can obtain new responses of the same statements but with better citations $R^{*}=\{r_{1},e^{*}_{1},...,r_{S},e^{*}_{S}\}$. Such preference data can be used in direct preference optimization (DPO) (Rafailov et al., 2024) to align the model based on the preference between the original outputs and improved outputs. Instead of using DPO, we choose its variant SimPO (Meng et al., 2024) here, as SimPO does not require a reference model and allows 2$\times$ memory saving for 25.6K long-context fine-tuning. Through this self-supervised process, which does not require ground-truth answers or human annotations, the model learns to generate more accurate and contextually grounded citations on its own.
@@ 3.1 Model Details
We use LongCite-8B, the Llama-3.1-8B model (Dubey et al., 2024) fine-tuned on LongCite-45K SFT data (Zhang et al., 2024) as the start point for both best-of-N sampling and preference optimization. We adopt the same text segmentation strategy from Zhang et al. (2024): each document is split into individual sentences using NLTK (Bird, 2006) and Chinese punctuations, and each sentence is prepended with a unique identifier in <C{$i$}> format. These identifiers serve as the citation indices, enabling the model to cite relevant context right after the statements with the format of <statement> {content ...} <cite>[$i_{1}-i_{2}$][$i_{3}-i_{4}$]...</cite></statement>. This format allows the model to cite a single sentence (e.g. $i_{1}=i_{2}$) or a span (e.g. $i_{1}<i_{2}$) efficiently within several tokens. The responses are generated via top-p sampling (Holtzman et al., 2020) with p=0.7 and temperature=0.95. We set p=0.9 and temperature=1.2 when doing best-of-N sampling for the citation strings to increase the diversity. We set N=10 in all the experiments considering the limited diversity in citations. [^1] After deduplicating repeated citation candidates, on average there are only 4.8 candidates left per statement in the BoN experiment on LongBench-Cite, with a standard deviation of 3.2.
@@ 3.2 Preference Optimization
Since best-of-N responses tend to have slightly longer citations, directly fine-tuning on them can lead the model to adopt a shortcut—generating longer citations instead of improving citation quality. To prevent this, we introduce length balancing: if an original response has a shorter citation length than the best-of-N response, we insert random citations from nearby sentences. This encourages the model to focus on where to cite rather than simply citing more. Details are provided in Appendix C, with an ablation study in Section 4.2.
Table 1: Citation recall (R), citation precision (P), citation F1 (F1), and citation length evaluated on LongBench-Cite benchmark. The best of our results are bolded. The best of previous state of the art are underlined. indicates the results taken from Zhang et al. (2024).
@@
3.3 Evaluation
Benchmark.
Baselines.
SelfCite is compared with these baselines.
  • Contributive context attribution: Contributive context attribution seeks to directly identify the parts of the context that cause the model to generate a particular statement. We consider ContextCite (Cohen-Wang et al., 2024), a contributive context attribution method that performs several random context ablations to model the effect of ablating different parts of the context on a generated statement. We use NLTK to split Llama-3.1-8B-Instruct’s responses into statements, and then apply ContextCite with 32 and 256 times of random context ablations to get the citations, with the details described in Appendix B.
  • Fine-tuned models: LongCite-8B and 9B released by Zhang et al. (2024), trained on LongCite-45K, fine-tuned from Llama-3.1-8B (Dubey et al., 2024) and GLM-4-9B (GLM et al., 2024), respectively. Additionally, we consider a baseline of finetuning LongCite-8B using SimPO with the NLI rewards which resembles Huang et al. (2024), with the details in Appendix E.
@@ 3.4 Main Results
Table 1 presents our main results. Our best-of-N sampling (BoN) consistently improves both citation recall and citation precision across tasks, increasing the overall F1 score from 73.8 to 77.5. Using SimPO to internalize BoN’s gains—eliminating the need for costly BoN sampling—achieves a similar improvement, with an F1 of 77.9. Applying BoN again to the SimPO fine-tuned model further boosts F1 by 5.3 points to 79.1, the highest across the datasets, suggesting room for further gains. Our results surpass LongCite-8B/9B at similar citation lengths and outperform proprietary model prompting while producing shorter citations.
To better contextualize the gains of our proposed reward, we additionally implement a variant of SimPO using NLI-based citation precision/recall rewards from Huang et al. (2024) by using the same training pipeline and initialization as our SimPO, modifying only the reward function (see details in Appendix E). As shown in row of SimPO w/ NLI Rewards, this baseline improves LongCite-8B on 3 out of 5 datasets, but is still consistently outperformed by SelfCite. This result highlights that while NLI-based rewards are helpful, our SelfCite reward provides a more accurate signal for optimizing citation quality.
Besides the fine-tuned baselines, we additionally compare our method to ContextCite for reference, a method very different from SelfCite–it does not directly generate citations, it estimates the importance scores of the context sentences after the response is generated (in Appendix B we show how to convert continuous importance scores into citations). Both SelfCite and ContextCite rely on the idea of context ablation, but our approach is significantly better. A key reason is that ContextCite estimates sentence importance from scratch using linear regression, while we rerank existing LLM-generated citation candidates, leading to more efficient and accurate citation quality estimation.
Finally, we evaluate the latest released Claude Citations API, as shown in Appendix D that SelfCite achieves strong results very close to this commercial-level API, validating the effectiveness of SelfCite.
@@
In our main experiment, we start from the Llama-3.1-8B model fine-tuned on the LongCite-45K SFT data, which effectively kick-starts its ability to generate structured citations for best-of-N sampling. The subsequent SimPO alignment stage is entirely self-supervised. We are also curious if it is possible to start from a fully self-supervised SFT model and then apply our self-supervised alignment after that. To begin with, we automatically generate 11K citation SFT data using ContextCite (see Appendix B for details) to replace the LongCite-45K annotations in the training data, as shown in the results at the bottom of Table 1. We can see that SFT on ContextCite can achieve decent initial results (65.7 F1) but still far from LongCite-8B (73.8 F1). BoN helps improving F1 to 67.3. After SimPO training, it achieves 69.9 F1, and additionally applying BoN can boost its F1 by 5.8 to 71.5, significantly closing the gap to LongCite-8B, showing our alignment method not only improve the supervised models, but also enhance the models purely trained from self-supervision.
@@
For best-of-N sampling, only the citation parts are modified, so the responses it generates to answer the questions are the same as those of the original LongCite-8B model, maintaining the same correctness. For the SimPO fine-tuned models, we test their answer correctness by the evaluation in Zhang et al. (2024), which contains two settings: answering with/without citations. If answering with citations, the model will be prompted to generate answers with structured citations, making the task more complex, and the citation parts will be removed when evaluating the answer correctness. The results in Table 2 show that the SimPO fine-tuning does not change the correctness of the LongCite-8B model much. The correctness is similar to LongSFT-8B/9B (Zhang et al., 2024), which are ablation baselines fine-tuned on LongCite-45k QA pairs but without the citation parts. The same observation still holds when starting from Llama-3.1-8B-Instruct, either SFT with ContextCite data or the further SimPO step do not change the answer correctness significantly. Under the same answer correctness, the additional “citations” can benefit the verifiability of the answers, enabling a user to easily double-check the answer, even in cases where the answers are wrong.
Chunk-level Citation Evaluation.
Additionally, we evaluate our methods on the traditional chunk-level citation benchmark ALCE (Gao et al., 2023b). However, due to the mismatch of data distributions and different task settings during training (sentence-level) and evaluation (chunk-level), we consider this as a zero-shot evaluation, and the results are shown in Appendix F, due to the limited space.
@@ 4.1 Ablation Study on Rewards
To better understand our final reward design, we explore various reward strategies in the BoN sampling process. Here, all BoN candidates are pre-generated and fixed, the reward is the only factor affecting results. Table 3 presents our ablation results on HotpotQA, while citation lengths are computed across all LongBench-Cite datasets for direct comparison with Table 1. We evaluate four alternative reward designs. BoN by LM log prob re-ranks candidates simply by the probability of the citation string, <cite>[$i_{1}-i_{2}$][$i_{3}-i_{4}$]...</cite>, which is similar to beam search but less costly. We observe that this strategy slightly boosts recall while reducing precision, resulting in a minor reduction in F1. BoN by max citation length always selects the candidates with the longest citations, i.e. citing the greatest number of sentences. Although it improves recall, it significantly reduces precision from 77.9 to 73.6 and inflates the citation length from 83.5 to 139.8. By contrast, both BoN by Prob-Drop and BoN by Prob-Hold improve recall without sacrificing precision. Finally, by combining both Prob-Drop and Prob-Hold into our final SelfCite reward, we achieve the best outcome, increasing both recall and precision and a 4-point improvement in F1.